There are n equal resistors. Series connection of these resistors gives 98 Ohms. Parallel connection of these resistors gives 2 Ohms. Find n.
1
Expert's answer
2020-12-17T09:01:06-0500

Solution:


R*n = 98


"\\dfrac{R}{n}\t=2" R/n = 2 


2n*n = 98


n = "\\sqrt{\\smash[b]{ \\tfrac{98}{2}\t}}" = 7 


R = 14 ohm 


check : 

14*7 = 70+28 = 98 ohm


"\tfrac{14}{7}=2 ohm"


Answer:

n=2